### v2

# FTFT: Efficient and Robust Fine-Tuning by Transferring Training Dynamics

Abstract
Despite the massive success of fine-tuning Pre-trained Language Models (PLMs), they remain susceptible to out-of-distribution input. Dataset cartography is a simple yet effective dual-model approach that improves the robustness of fine-tuned PLMs. It involves fine-tuning a model on the original training set (i.e. reference model), selecting a subset of important training instances based on the training dynamics, and fine-tuning again only on these selected examples (i.e. main model). However, this approach requires fine-tuning the same model twice, which is computationally expensive for large PLMs. In this paper, we show that 1) training dynamics are highly transferable across model sizes and pre-training methods, and that 2) fine-tuning main models using these selected training instances achieves higher training efficiency than empirical risk minimization (ERM). Building on these observations, we propose a novel fine-tuning approach: Fine-Tuning by transFerring Training dynamics (FTFT). Compared with dataset cartography, FTFT uses more efficient reference models and aggressive early stopping. FTFT achieves robustness improvements over ERM while lowering the training cost by up to .111 Our code is publicly available at https://github.com/nlpsoc/FTFT.
1 Introduction
Despite the success of few-shot and zero-shot learning (Brown et al., 2020), state-of-the-art performance in Natural Language Processing (NLP) is still largely achieved by fine-tuning large Pre-trained Language Models (PLMs) (Mosbach et al., 2023). Scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022) suggest that better downstream performance is achieved with larger PLMs. However, fine-tuning large PLMs is also more expensive, in terms of both computational resources and carbon emission (Strubell et al., 2019; Wu et al., 2022).
Moreover, despite impressive progress on regular benchmarks, many studies have shown that fine-tuned PLMs lack robustness against out-of-distribution (OOD) input. For instance, human annotators can easily exploit the weaknesses of fine-tuned PLMs to trick these models to yield incorrect predictions, on tasks such as Natural Language Inference (NLI) (Nie et al., 2020) and Hate Speech Detection (HSD) (Vidgen et al., 2021b).
The problem of robustness can be mitigated using dual-model approaches. With such approaches, first a reference model is trained to estimate the importance of each training instance, and then a main model is trained based on the outputs of the reference model (Nam et al., 2020; Utama et al., 2020; Sanh et al., 2021; Karimi Mahabadi et al., 2020; Zhang et al., 2022; Liu et al., 2021). Of these approaches, dataset cartography (Swayamdipta et al., 2020) is especially attractive in view of its simplicity and the consistent improvements in robustness that it yields. It consists of three steps. First, a Data Map (DM) is constructed from the training dynamics (i.e. instance prediction probabilities) of a fine-tuning run of the reference model on the full dataset. This DM divides the training data into three subsets: ambiguous, hard-to-learn, and easy instances. Finally, the main model is fine-tuned using only either the ambiguous or hard-to-learn subset. An important question has to do with the choice of reference model. Swayamdipta et al. (2020) use the same PLM for both the reference and main model. However, a major drawback is the high computational cost, because it requires fine-tuning the same model twice. In this paper, we jointly address robustness and efficiency issues without sacrificing the simplicity of dataset cartography, by exploiting the transferability of training dynamics. We make three contributions.
First, we study the following question: Are DMs transferable across different model sizes and pretraining methods? We focus on the novel setting where DMs are constructed based on computationally efficient reference models to fine-tune more capable — and often larger — main models. The motivation is two-fold: 1) efficient reference models can ease the computational burden of constructing DMs, and 2) less capable reference models could be better at identifying ambiguous or hard training instances, because they are less likely to memorize training data (Tirumala et al., 2022; Carlini et al., 2023). We observe that, in most cases, training dynamics are highly transferable across different model sizes (§4.1) and pretraining methods (§4.2). We further show that the condition for successful transfers is a reasonably strong reference model, in a sense that we also seek to make precise (§4.3).
Second, we observe that fine-tuning using training instances selected by DMs enjoys consistently higher training efficiency than conventional fine-tuning (§6).
Third, building on these findings, we propose Fine-Tuning by transFerring Training dynamics (FTFT, §6): an efficient fine-tuning approach that leads to improved OOD performance. Compared to dataset cartography, FTFT uses more efficient reference models and early stopping. Experiments on two tasks, NLI and HSD, show that FTFT achieves better performance on OOD input than a standard approach based on Empirical Risk Minimization (ERM), while lowering the training cost by up to .
2 Background
Dual-Model Approaches for Robustness
Many studies have proposed dual-model approaches to improve model robustness that do not require knowledge of identifiable subsets of the data samples, such as NLI pairs in which hypotheses contain a negation (Gururangan et al., 2018), or HSD samples targetting a specific group (Dixon et al., 2018; Park et al., 2018). Nam et al. (2020) first train a reference model using generalized cross-entropy loss, and then train a main model while assigning higher weights to instances that are hard for the reference model. Sanh et al. (2021) use a Product-of-Expert (PoE) approach, by first training a reference model with limited capacity to capture dataset biases, and then training the main model to avoid these biases using PoE loss. Liu et al. (2021) propose the Just-Train-Twice approach (JTT), which involves first training a weak reference model using heavy regularization and vanilla SGD, and then up-weighing the training instances that the reference model predicts incorrectly when training the main model. Dataset cartography (Swayamdipta et al., 2020) is based on a similar idea, but it uses training dynamics instead of correctness to categorize training instances. We discuss this method below.
Dataset Cartography
is a dual-model approach to improve model robustness. First, a reference model is trained on the full training dataset. Then, a Data Map (DM) is built based on the training dynamics, by tracking prediction probabilities of the true class () of each training instance across epochs. DM categorizes training instances into three subsets: ambiguous (i.e. the variance of is in the top of all training instances); hard-to-learn (i.e. the mean of is in the bottom of all training instances); and easy (i.e. neither ambiguous nor hard-to-learn). The threshold is fixed and typically set to . Note that a training instance can be categorized as both hard-to-learn and ambiguous (a low mean but high variance for ). Finally, the main model is fine-tuned only on the ambiguous or hard-to-learn subset. Swayamdipta et al. (2020) show that, with a slight loss of In-Distribution (ID) performance, dataset cartography improves Out-Of-Distribution (OOD) performance of models. In this paper, we train main models with ambiguous data, since Swayamdipta et al. (2020) reported better performance using this data than hard-to-learn data.
Swayamdipta et al. (2020) use the same PLM as both the reference and the main model. In contrast, Sar-Shalom & Schwartz (2023) show that a DM constructed by (Clark et al., 2020) can be used to improve the robustness of (He et al., 2023). However, instead of using only the ambiguous subset, they added copies of this subset to the original training set to train the main model. Moreover, they did not investigate either DM transfer across model sizes and pretraining methods, or how such transferability can be exploited to improve efficiency.
Model-Based Data Selection/Reweighing
Our work is also connected to studies that have selected or reweighed data using reference models to improve in-distribution (ID) performance. Chang et al. (2017) use variance and proximity to the classification threshold from a reference model to reweigh training instances; Toneva et al. (2019) calculate the frequency of forgetting events (i.e. from correct to incorrect prediction), and remove the least forgettable instances; Paul et al. (2021) and Baldock et al. (2021) instead use error vector norm and effective prediction depth to estimate the contribution of a training instance.
Previous studies have also explored the use of a smaller reference model to improve efficiency. Coleman et al. (2020) use a small model for active learning and core-set selection. Xie et al. (2023) reweigh domains for language model pretraining, by training a small reference model to estimate the difficulty of each domain.
3 Experimental Setup
We perform our experiments on two tasks, Natural Language Inference (NLI) and Hate Speech Detection (HSD). Following Swayamdipta et al. (2020), we set .
Data
To study model robustness, we include a few challenging OOD test sets for each task, in addition to the training set and an ID validation set. For NLI, we use the MultiNLI dataset (Williams et al., 2018) as the train and ID validation set, because of its diverse composition, covering 10 genres. We use AdversarialNLI (Nie et al., 2020) as the OOD test set, which consists of three rounds of adversarial data collection. AdversarialNLI is known to be challenging and it targets weaknesses of models trained on MultiNLI.
For HSD, we use CAD (Vidgen et al., 2021a) as the training and ID validation set. CAD consists of Reddit posts covering diverse topics and writing styles, annotated based on a fine-grained taxonomy. Following Ramponi & Tonelli (2022), we frame the task as a binary classification task, by marking identity-related abuse as hateful and other categories as non-hateful. As OOD test sets, we use DynaHate (Vidgen et al., 2021b), which contains three rounds of adversarial data collection and perturbations, and aligns with CAD’s hate speech definition (i.e. the datasets have the same labeling scheme).
Models
We mainly use DeBERTaV3 (He et al., 2023) and ELECTRA (Clark et al., 2020) in our experiments, because they offer strong performance and various model sizes. To study the transferability across different pretraining methods, we also use TinyBERT (Turc et al., 2020), BERT (Devlin et al., 2019) and RoBERTa (Liu et al., 2019) as reference models.222 Costs for fine-tuning different PLMs are in Appendix A.1. We report FLOPs rather than GPU hours because we noticed occasional low GPU utilization especially when fine-tuning smaller PLMs. Besides ERM, we include a baseline that uses a random DM (i.e., random of the train data). We perform a search of optimal training steps on ERM, and use the same hyper-parameters for both ERM and data cartography approaches. Full details of the models and training setups are in Appendix A.1.
4 Transferability of Training Dynamics
In this section, we study the transferability of training dynamics in dataset cartography, i.e., whether we can use different reference and main models while maintaining the robustness advantage of the main model. Specifically, we study whether training dynamics are transferable across different model sizes (§4.1, e.g., from to ) and pretraining methods (§4.2, e.g., from to ).
We focus on these issues for two reasons. First, transferability across model sizes enables using more efficient (and usually less capable) reference models, which (1) can improve training efficiency, and (2) is potentially more effective in identifying ambiguous/hard instances, because they are usually worse at memorizing training instances. Second, transferability across pretraining methods can help achieve these advantages even in cases where more efficient variants of the main pretraining method are unsuitable or unavailable for the task. Moreover, understanding transferability can shed light on data importance: If DMs of different reference models consistently identify the same subset of training instances as ambiguous, it suggests that DMs reveal intrinsic data characteristics.
We define successful transfers as transfers that produce comparable or better OOD performance than ERM. Our results show that, with a few exceptions, training dynamics are indeed transferable. To understand the conditions for successful transfers, we analyze the failure cases (§4.3), and find that the DMs of reference models that lead to successful transfers typically identify a larger subset as easy, which serves as a rough indicator of their capability. This finding can serve as a guideline for choosing reference models without training the main model, which is computationally expensive.
4.1 Transferability Across Model Sizes
In this section, we study whether smaller and more efficient models can be used as reference models to construct DMs for training larger main models (e.g. as the reference model for ). Levering transferability of this type can improve the efficiency of dataset cartography by reducing the cost of constructing DMs.
The results for DeBERTaV3 are shown in Table 1 under section across different model sizes (NLI) and Table 5 (HSD, Appendix B). Training dynamics are transferable across different model sizes: when using as the main model, changing the reference model to either or yields comparable or even better performance. This observation is consistent with our hypothesis that efficient models are more sensitive to ambiguous and difficult instances. As a result, they can serve as more effective reference models compared to larger, more capable models. We also observe that, consistent with Swayamdipta et al. (2020), ERM achieves better ID performance (in the baselines section of Table 1), while data cartography performs comparably or better on OOD data. Note that hyper-parameters are determined by ERM performance (see Appendix A.1), making the training setup more favorable overall to ERM.
To investigate transferability of DMs further, we analyze whether reference models of different sizes identify similar groups of ambiguous instances. Figure 1(a) shows the percentage of ambiguous instances shared by reference models of different sizes and random seeds. We observe that the percentages of shared ambiguous instances between different sizes are only slightly smaller than those between the same size but different random seeds, providing further evidence for the transferability of DMs.333 The expected overlap between two random DMs is 0.33.
4.2 Transferability Across Pretraining Methods
Method Main Model Ref. Model Cost MultiNLI AdversarialNLI (Test) - R1 R2 R3 Baselines: ERM, ERM with Early Stopping, and Random DM ERM - 14.47 ERM - 28.29 ERM* - 100.00 ERM(ES) - 66.67 DM Random 100.00 Training Dynamics Transferability: Across Different Model Sizes DM 200.00 DM 114.47 DM 128.29 Training Dynamics Transferability: Across Different Pretraining Methods DM 104.61 DM 136.18 DM 216.78 DM 213.49 FTFT: Efficient Reference Models + Aggressive Early Stopping FTFT 51.97 FTFT 74.12 FTFT 79.93
We now study the transferability of training dynamics across different pretraining methods. Successful transfers of this type enable the exploitation of efficient reference models, in cases where there is no efficient version of the main model that suits the downstream task.
The results for as the main model with different reference models are shown in Table 1 under section across different pretraining methods (NLI) and Table 5 (HSD, Appendix B). We observe that training dynamics are generally transferable across different pretraining methods: achieves comparable performance using DMs constructed by different reference models in most cases. However, there are a few exceptions: when using as the reference model, the performance is clearly worse on the NLI OOD datasets than when using ERM. We hypothesize that is not strong enough for constructing effective DMs, and analyze this hypothesis further in §4.3 below.
4.3 How Efficient Can We Be?
Method Main Model Ref. Model Compute MultiNLI (Val.) AdversarialNLI (Test) - R1 R2 R3 ERM TinyBERT - 1.45 ERM - 4.61 ERM - 36.18 ERM - 14.47 ERM - 28.29 ERM - 113.49 ERM - 116.78 DM TinyBERT 101.45 DM 104.61 DM 136.18 DM 114.47 DM 128.29 DM 213.49 DM 216.78 DM TinyBERT 111.64 DM 114.80 DM 146.38 DM 124.67 DM 138.49 DM 223.68 DM 226.97
We have shown that training dynamics are usually transferable across different model sizes and pretraining methods. We now study the conditions for successful transfers, by zooming in on two questions: 1) Can we use efficient but weak models as reference models? and 2) What are the differences between effective and ineffective reference models? Answers to these questions can guide the selection of efficient yet effective reference models.
Can we use efficient but weak models as reference models?
To answer this question, we compare the performance of a wide range of methods, see Table 2 (NLI) and Table 6 (HSD, in Appendix B). We include models from three categories: First, a wide range of seven models trained using ERM: TinyBERT, the small and base versions of DeBERTaV3 and ELECTRA, , and . Second, we use these ERM models as reference models to construct DMs, and use these DMs to fine-tune . By using reference models with different sizes and pretraining methods to fine-tune the same main model, we can inspect the impact of reference model capability on transferability. Third, we also include the results of using as the main model. By comparing results from different main models, we can better understand whether successful transfers originate from the compatibility between reference and main models, or the capability of the reference model itself. We make three observations.
First, weak reference models with poor ID performance, i.e., TinyBERT and , lead to unsuccessful transfers. However, the reference models do not need to be as capable as the main model: slightly weaker reference models could be more useful, indicated by the strong OOD performance of using as the reference model. Moreover, in these unsuccessful transfers, the main model OOD performance correlates with the reference model ID performance. For example, on NLI, TinyBERT has a lower ID performance than , and both main models and have a lower OOD performance when using TinyBERT as the reference model compared to when using as the reference model.
Second, whether a transfer is successful or not depends mostly on the reference model, rather than the compatibility between the reference and the main models: The transfers from TinyBERT and to both main models are unsuccessful, and the transfers from other reference models to both main models are successful.
Third, interestingly, ID performance in all transfers remains high. In other words, ineffective reference models only impact the OOD performance of the main models. For instance, the transfer from to yields the best accuracy on MultiNLI despite its OOD performance being relatively poor. We suspect that weak models with poor ID performance often identify easy training instances as ambiguous data. While easy instances can be sufficient for obtaining satisfactory ID performance, they are insufficient for good OOD performance (Swayamdipta et al., 2020). We discuss this in detail below.
What are the differences between effective and ineffective reference models?
To answer this question, we consider the possible differences between a weak (and ineffective) reference model and a reasonably strong reference model, in terms of categorizing training data into ambiguous, hard-to-learn, and easy subsets. Also, we assume that instances in our training set exhibit varying levels of difficulty (i.e. simple to difficult).
Assume we have a weak reference model that can learn simpler training instances but cannot learn the more difficult ones. This weak reference model will therefore assign increasing to simple training instances across different epochs, while keeping for difficult training instances around the values expected in a random guessing scenario. Consequently, will exhibit high standard deviations on simpler training instances, which will then be identified as ambiguous data; while more difficult training instances will have consistent lower mean values and thus low standard deviations for , and therefore be identified as hard-to-learn data. In contrast, a reasonably capable reference model can learn simpler training instances during the early stage of training (even before their first occurrence in the training batch, i.e., their correct predictions are learned from other training instances). Therefore, these instances will have both high mean values and low standard deviations for . Meanwhile, for difficult instances will gradually increase across epochs, making these instances yield relatively low mean values and high standard deviations for . As a result, these instances will be identified as both ambiguous and hard-to-learn (i.e. we expect a large overlap in these subsets). Because we select a fixed percentage of instances as ambiguous or hard-to-learn, this larger overlap means a larger easy subset too.
We now validate our reasoning. Given a reference model, we first split the training instances into two subsets based on mean : hard-to-learn ( of training instances) and other (the remaining ). We use a lower to make the difference clearer. Then, for each subset, we calculate the median in each epoch. We use median values because they are robust statistics of the central tendency. Figure 1(b) shows our results on MultiNLI, using two effective ( and ) and two ineffective ( and TinyBERT) reference models. We only include four models to make the figure clearer — we observe similar trends for other models: With effective reference models, hard-to-learn instances are gradually learned during training, while other instances already have high values from the first epoch. In contrast, with ineffective reference models, hard-to-learn data instances are not learned at all, suggested by their close-to-zero over different epochs; while other instances are gradually fitted, indicated by their increasing values.
To further validate our reasoning, we compute the percentages of training instances identified as easy by different reference models (Table 4 in Appendix B): We report results using to obtain more comprehensive results. Less effective reference models indeed identify fewer data points as easy. For example, on NLI with , TinyBERT identifies less than 20.0% of the instances as easy, compared to 46.65% by . Furthermore, the overlap between hard-to-learn and ambiguous instances in successful transfers is usually very large. For example, with , all effective reference models identify more than 46% as easy training instances (the maximum is 50%, when ambiguous and hard-to-learn data align perfectly).
5 Training Speed Gains From Data Selection
Our results from §4 suggest that one can improve the efficiency of dataset cartography with more efficient reference models. However, if we train the main models for the same number of steps as ERM, the cost of the full pipeline is still higher than conventional fine-tuning, because of the extra reference model training cost.
Previous studies have shown that training with a careful selection of “informative” training instances can lead to higher training speed than ERM (i.e. achieving the same performance with fewer training steps), despite the computational cost of selecting these instances being notoriously high (Sorscher et al., 2022; Feldman & Zhang, 2020; Paul et al., 2021; Toneva et al., 2019). Motivated by this, we now study whether we can achieve similar learning speed gains by training on instances selected by DMs. If such gains exist, we can further improve the efficiency of data cartography by training with fewer steps.
We present the performance of on HSD, fine-tuned using ERM and training instances selected by DMs, across different training durations (i.e. max training steps) in Figure 2. We include results for other datasets in Appendix B, on which we observe similar trends. Specifically, we make three observations. First, when the number of training steps is reduced, models fine-tuned with dataset cartography clearly outperform those fine-tuned with ERM, even in cases where ERM achieves better performance with more training steps (i.e., on ID validation data). Second, dataset cartography often yields better results with fewer training steps than with more, suggesting a tendency of overfitting in the late stage of training. Third, such efficiency gains hold consistently across different training durations. All three observations indicate that training with data instances selected by DMs achieves consistent higher training speed than ERM.
6 Our Approach: FTFT
Building on our insights presented in the previous sections, we now propose a novel fine-tuning approach based on dataset cartography, Fine-Tuning by transFerring Training dynamics (FTFT). Compared with the original dataset cartography approach, FTFT integrates two crucial changes: (1) using more efficient reference models, which not only result in comparable DMs but also enhance efficiency, and (2) implementing aggressive early stopping during the main model training. Such aggressive early stopping is possible because models trained with data instances selected using DMs already achieve strong performance with substantially fewer training steps (§5).
We show our results on NLI in Table 1 (FTFT) and on HSD in Table 5 (Appendix B). We use an early stopping patience (i.e. the number of checkpoints without improvement before stopping), based on the average dev set performance of the OOD datasets. We use a relatively small , because it can stop the training earlier, leading to higher efficiency. We also show the computational cost of each method (Compute), in which we have taken the training of both the reference model and the main model into account, including the extra checkpoints after the best performance is achieved.
We make three observations. First, FTFT achieves consistent robustness improvement over ERM, indicated by its strong performance on most OOD datasets (note again that the hyper-parameters are optimized for ERM). We also include the results of ERM with early stopping (ERM(ES)) as a baseline to illustrate that our aggressive early stopping strategy is only effective when training with data instances selected by DMs: ERM(ES) achieves worse performance than ERM on all OOD datasets. Second, FTFT achieves substantial higher training efficiency than other methods. For example, on NLI, training using FTFT and as reference models only costs 51.57% of the training cost of ERM, which is equivalent to 25.78% of the training cost of the original dataset cartography approach. Third, our early stopping strategy is effectively aggressive. Specifically, FTFT with all three reference models in Table 1 achieves their best performance within just 1/3 of the optimal training steps of ERM.
7 Conclusions
Fine-tuned PLMs have been shown to be vulnerable to OOD inputs. Although dataset cartography can improve model robustness (Swayamdipta et al., 2020; Sar-Shalom & Schwartz, 2023), it is computationally expensive. In this paper, we have presented FTFT, a novel approach for fine-tuning PLMs which yields both better efficiency and better robustness over ERM (§6). FTFT is built on dataset cartography, based on two observations: (1) reference model training dynamics are highly transferable across different model sizes (§4.1) and pretraining methods (§4.2), and (2) main models trained using data instances selected by reference model training dynamics achieve higher learning speed (i.e. substantially better performance with reduced number of training steps). We have also discussed the conditions for successful FTFT runs (§4.3). We believe that FTFT will be an important tool for future researchers and practitioners to perform efficient model fine-tuning, especially in situations where robustness is essential. We also believe our study of lightweight reference models will benefit future research on building efficient supervised data pruning metrics, which are shown to outperform their unsupervised counterparts (Sorscher et al., 2022).
Our work opens up several directions for future research. First, we have observed that effective reference models identify more instances as easy. Further controlled experiments are needed to build a more rigorous protocol for choosing reference models. We believe the results presented here provide a good foundation for such further work. Second, we have empirically identified the transferability of training dynamics to selected training instances. Future studies are needed to build the theoretical foundations of both data cartography itself and the feasibility of such transfers. Third, we have only developed FTFT for classification tasks. Future studies can extend our work to other tasks, such as generation (e.g., question answering) and self-supervised tasks (e.g., language modeling).
Appendix A Training Specifications
A.1 Experimental Setup
Optimization
For training all models, we use AdamW (Loshchilov & Hutter, 2019) as the optimizer with a batch size of 32. We also use a linear learning rate scheduler with 10% warmup. For fine-tuning the small and base versions of both DeBERTaV3 and ELECTRA, as well as TinyBERT, RoBERTa, and BERT, we use a learning rate of 2e-5, following the suggestions from the original papers. We use a smaller learning rate of 1e-5 for and , because training larger models with lower learning rates are observed to be more stable (Mosbach et al., 2021; Du & Nguyen, 2023). However, a few failed training runs (i.e., training fails to converge and the resulting model performs worse than the majority class baseline Mosbach et al., 2021) are still observed during the training of . We excluded these runs from our results.
Number of Training Steps and Checkpointing
Regarding the number of training steps, we perform a grid search for both and ERM models, spanning from one to five epochs, according to their average performance on the validation set of AdversarialNLI (NLI) and DynaHate (HSD). On NLI, the optimal length of training is four epochs (49088 steps) and five epochs (61360 steps) for and . On HSD, the optimal length of training is three epochs (1620 steps) for both and . For other PLMs, because we only use them as reference models, we do not perform this grid search, and use 61360 steps and 1620 steps for NLI and HSD. We perform checkpointing every 4090 steps and 180 steps on NLI and HSD, which is approximately the length of one epoch using the 33% selected data instances from DMs.
Software and Hardware
We use Python 3.9 and PyTorch 2.0 for all experiments. We also use HuggingFace Transformers 4.32 (Wolf et al., 2020), Accelerate 0.22, and Datasets 2.14 (Lhoest et al., 2021). All experiments are performed on one NVIDIA A100 GPU. Training all models (including hyper-parameter search) takes approximately 16 GPU days.
A.2 Comparison of Training Costs
| Model | Param Size | Compute |
| 44.00M | 14.47 | |
| 86.00M | 28.29 | |
| 304.00M | 100.00 | |
| 14.00M | 4.61 | |
| 110.00M | 36.18 | |
| 335.00M | 110.20 | |
| 345.00M | 113.49 | |
| 355.00M | 116.78 | |
| 4.40M | 1.45 |
Appendix B Additional Results
NLI: MultiNLI HSD: CAD 10% 25% 33% 50% 10% 25% 33% 50% Model 80.01% 51.27% 38.34% 19.60% 81.89% 65.48% 58.20% 41.73% 80.10% 55.72% 47.40% 35.50% 82.97% 69.70% 61.92% 41.51% 84.57% 69.00% 62.84% 46.65% 86.10% 72.07% 63.41% 44.54% 82.80% 67.54% 61.72% 46.15% 85.44% 70.82% 62.83% 46.54% 85.08% 71.58% 64.06% 46.86% 85.49% 72.30% 63.64% 46.14% 85.21% 70.08% 63.05% 47.21% 89.25% 72.38% 64.15% 47.07% 85.67% 71.43% 64.24% 47.20% 88.57% 73.21% 64.84% 47.25%
| Method | Main Model | Ref. Model | Compute | CAD (Val.) | DynaHate (Test) | |||||
| - | Ori-R2 | Ori-R3 | Ori-R4 | Pert-R2 | Pert-R3 | Pert-R4 | ||||
| Baselines: ERM, ERM with Early Stopping, and Random DM | ||||||||||
| ERM | - | 14.47 | ||||||||
| ERM | - | 28.29 | ||||||||
| ERM* | - | 100.00 | ||||||||
| ERM(ES) | - | 69.44 | ||||||||
| DM | Random | 100.00 | ||||||||
| Training Dynamics Transferability: Across Different Model Sizes | ||||||||||
| DM | 200.00 | |||||||||
| DM | 114.47 | |||||||||
| DM | 128.29 | |||||||||
| Training Dynamics Transferability: Across Different Pretraining Methods | ||||||||||
| DM | 104.61 | |||||||||
| DM | 136.18 | |||||||||
| DM | 216.78 | |||||||||
| DM | 213.49 | |||||||||
| FTFT: Efficient Reference Models + Aggressive Early Stopping | ||||||||||
| FTFT | 78.36 | |||||||||
| FTFT | 106.07 | |||||||||
| FTFT | 108.40 | |||||||||
| Method | Main Model | Ref. Model | Compute | CAD (Val.) | DynaHate (Test) | |||||
| - | Ori-R2 | Ori-R3 | Ori-R4 | Pert-R2 | Pert-R3 | Pert-R4 | ||||
| ERM | TinyBERT | - | 1.45 | |||||||
| ERM | - | 4.61 | ||||||||
| ERM | - | 36.18 | ||||||||
| ERM | - | 14.47 | ||||||||
| ERM | - | 28.29 | ||||||||
| ERM | - | 113.49 | ||||||||
| ERM | - | 116.78 | ||||||||
| DM | TinyBERT | 101.45 | ||||||||
| DM | 104.61 | |||||||||
| DM | 136.18 | |||||||||
| DM | 114.47 | |||||||||
| DM | 128.29 | |||||||||
| DM | 213.49 | |||||||||
| DM | 216.78 | |||||||||
| DM | TinyBERT | 111.64 | ||||||||
| DM | 114.80 | |||||||||
| DM | 146.38 | |||||||||
| DM | 124.67 | |||||||||
| DM | 138.49 | |||||||||
| DM | 223.68 | |||||||||
| DM | 226.97 | |||||||||

### v3

@@
 FTFT achieves robustness improvements over ERM while lowering the training cost by up to .111 Our code is publicly available at https://github.com/nlpsoc/FTFT.
+FTFT: Efficient and Robust Fine-Tuning by Transferring Training Dynamics
+Yupei Du, Albert Gatt, and Dong Nguyen Utrecht University Utrecht, the Netherlands {y.du, a.gatt, d.p.nguyen}@uu.nl
 1 Introduction
@@
 With such approaches, first a reference model is trained to estimate the importance of each training instance, and then a main model is trained based on the outputs of the reference model (Nam et al., 2020; Utama et al., 2020; Sanh et al., 2021; Karimi Mahabadi et al., 2020; Zhang et al., 2022; Liu et al., 2021).
-Of these approaches, dataset cartography (Swayamdipta et al., 2020) is especially attractive in view of its simplicity and the consistent improvements in robustness that it yields.
+Among these approaches, dataset cartography (Swayamdipta et al., 2020) is especially attractive in view of its simplicity and the consistent improvements in model robustness.
 It consists of three steps.
@@
 (2020) use the same PLM for both the reference and main model.
-However, a major drawback is the high computational cost, because it requires fine-tuning the same model twice.
+However, a major drawback of dataset cartography is the high computational cost, because it requires fine-tuning the same model twice.
 In this paper, we jointly address robustness and efficiency issues without sacrificing the simplicity of dataset cartography, by exploiting the transferability of training dynamics.
@@
 We focus on the novel setting where DMs are constructed based on computationally efficient reference models to fine-tune more capable — and often larger — main models.
-The motivation is two-fold: 1) efficient reference models can ease the computational burden of constructing DMs, and 2) less capable reference models could be better at identifying ambiguous or hard training instances, because they are less likely to memorize training data (Tirumala et al., 2022; Carlini et al., 2023).
-We observe that, in most cases, training dynamics are highly transferable across different model sizes (§4.1) and pretraining methods (§4.2).
-We further show that the condition for successful transfers is a reasonably strong reference model, in a sense that we also seek to make precise (§4.3).
-Second, we observe that fine-tuning using training instances selected by DMs enjoys consistently higher training efficiency than conventional fine-tuning (§6).
+Our motivation is two-fold: 1) efficient reference models ease the computational burden of constructing DMs, and 2) less capable reference models might be better at identifying ambiguous or hard training instances, because they are less likely to memorize training data (Tirumala et al., 2022; Carlini et al., 2023).
+Our results show that, in most cases, training dynamics are highly transferable across different model sizes (§4.1) and pretraining methods (§4.2).
+We further show that the condition for successful transfers is a reasonably strong reference model, which we also make precise in §4.3.
+Second, we observe that fine-tuning with selected instances achieves consistently higher training efficiency than conventional fine-tuning (§6).
 Third, building on these findings, we propose Fine-Tuning by transFerring Training dynamics (FTFT, §6): an efficient fine-tuning approach that leads to improved OOD performance.
-Compared to dataset cartography, FTFT uses more efficient reference models and early stopping.
-Experiments on two tasks, NLI and HSD, show that FTFT achieves better performance on OOD input than a standard approach based on Empirical Risk Minimization (ERM), while lowering the training cost by up to .
+Compared to dataset cartography, FTFT uses efficient reference models and early stopping.
+Experiments on two tasks, NLI and HSD, show that FTFT achieves better performance on OOD input than conventional Empirical Risk Minimization (ERM), while lowering the training cost by up to .
 2 Background
@@
 Nam et al.
-(2020) first train a reference model using generalized cross-entropy loss, and then train a main model while assigning higher weights to instances that are hard for the reference model.
+(2020) first train a reference model using generalized cross-entropy loss, and then train a main model while assigning higher weights to instances that were hard for the reference model.
 Sanh et al.
@@
 (2020) use the same PLM as both the reference and the main model.
-In contrast, Sar-Shalom & Schwartz (2023) show that a DM constructed by (Clark et al., 2020) can be used to improve the robustness of (He et al., 2023).
+In contrast, Sar-Shalom and Schwartz (2023) show that a DM constructed by (Clark et al., 2020) can be used to improve the robustness of (He et al., 2023).
 However, instead of using only the ambiguous subset, they added copies of this subset to the original training set to train the main model.
@@
 (2019) calculate the frequency of forgetting events (i.e.
-from correct to incorrect prediction), and remove the least forgettable instances; Paul et al.
+from a correct to incorrect prediction), and remove the least forgettable instances; Paul et al.
 (2021) and Baldock et al.
@@
 3 Experimental Setup
-We perform our experiments on two tasks, Natural Language Inference (NLI) and Hate Speech Detection (HSD). Following Swayamdipta et al. (2020), we set .
+We perform our experiments on two tasks, Natural Language Inference (NLI) and Hate Speech Detection (HSD). Following Swayamdipta et al. (2020), we select 33% the most ambiguous datapoints.
 Data
-To study model robustness, we include a few challenging OOD test sets for each task, in addition to the training set and an ID validation set.
+To study model robustness, we include challenging OOD test sets for each task, in addition to the training set and an ID validation set.
 For NLI, we use the MultiNLI dataset (Williams et al., 2018) as the train and ID validation set, because of its diverse composition, covering 10 genres.
@@
 For HSD, we use CAD (Vidgen et al., 2021a) as the training and ID validation set.
-CAD consists of Reddit posts covering diverse topics and writing styles, annotated based on a fine-grained taxonomy.
-Following Ramponi & Tonelli (2022), we frame the task as a binary classification task, by marking identity-related abuse as hateful and other categories as non-hateful.
-As OOD test sets, we use DynaHate (Vidgen et al., 2021b), which contains three rounds of adversarial data collection and perturbations, and aligns with CAD’s hate speech definition (i.e.
-the datasets have the same labeling scheme).
+CAD consists of Reddit posts covering diverse topics and writing styles, annotated using a fine-grained taxonomy.
+Following Ramponi and Tonelli (2022), we frame it as a binary classification task, marking identity-related abuse as hateful and other categories as non-hateful.
+As OOD test sets, we use DynaHate (Vidgen et al., 2021b), since it aligns with CAD’s definition of hate speech.
+DynaHate contains three rounds of adversarial data collection and perturbations.
 Models
-We mainly use DeBERTaV3 (He et al., 2023) and ELECTRA (Clark et al., 2020) in our experiments, because they offer strong performance and various model sizes.
+We mainly use DeBERTaV3 (He et al., 2023) and ELECTRA (Clark et al., 2020) in our experiments, due to their strong performance and availability in multiple model sizes.
 To study the transferability across different pretraining methods, we also use TinyBERT (Turc et al., 2020), BERT (Devlin et al., 2019) and RoBERTa (Liu et al., 2019) as reference models.222 Costs for fine-tuning different PLMs are in Appendix A.1.
 We report FLOPs rather than GPU hours because we noticed occasional low GPU utilization especially when fine-tuning smaller PLMs.
-Besides ERM, we include a baseline that uses a random DM (i.e., random of the train data).
+Besides ERM, we include a baseline that uses a random DM (i.e., random of the training data).
 We perform a search of optimal training steps on ERM, and use the same hyper-parameters for both ERM and data cartography approaches.
 Full details of the models and training setups are in Appendix A.1.
+Method Main Model Ref.
+Model Cost MultiNLI AdversarialNLI (Test) - R1 R2 R3 Baselines: ERM, ERM with Early Stopping, and Random DM ERM - 14.47 ERM - 28.29 ERM - 100.00 ERM(ES) - 66.67 DM Random 100.00 Training Dynamics Transferability: Across Different Model Sizes DM 200.00 DM 114.47 DM 128.29 Training Dynamics Transferability: Across Different Pretraining Methods DM 104.61 DM 136.18 DM 216.78 DM 213.49 FTFT: Efficient Reference Models + Aggressive Early Stopping FTFT 51.97 FTFT 74.12 FTFT 79.93
 4 Transferability of Training Dynamics
@@
 Our results show that, with a few exceptions, training dynamics are indeed transferable.
-To understand the conditions for successful transfers, we analyze the failure cases (§4.3), and find that the DMs of reference models that lead to successful transfers typically identify a larger subset as easy, which serves as a rough indicator of their capability.
+To understand the conditions for successful transfers, we analyze the failure cases (§4.3).
+We find that the DMs of reference models that lead to successful transfers typically identify a larger subset as easy, which serves as a rough indicator of their capability.
 This finding can serve as a guideline for choosing reference models without training the main model, which is computationally expensive.
@@
 This observation is consistent with our hypothesis that efficient models are more sensitive to ambiguous and difficult instances.
-As a result, they can serve as more effective reference models compared to larger, more capable models.
+As a result, they can serve as alternative efficient reference models.
 We also observe that, consistent with Swayamdipta et al.
 (2020), ERM achieves better ID performance (in the baselines section of Table 1), while data cartography performs comparably or better on OOD data.
-Note that hyper-parameters are determined by ERM performance (see Appendix A.1), making the training setup more favorable overall to ERM.
+Note that hyper-parameters are tuned for ERM performance (see Appendix A.1), making the training setup more favorable overall to ERM.
 To investigate transferability of DMs further, we analyze whether reference models of different sizes identify similar groups of ambiguous instances.
 Figure 1(a) shows the percentage of ambiguous instances shared by reference models of different sizes and random seeds.
-We observe that the percentages of shared ambiguous instances between different sizes are only slightly smaller than those between the same size but different random seeds, providing further evidence for the transferability of DMs.333 The expected overlap between two random DMs is 0.33.
+The percentages shared between different model sizes are only slightly smaller than those between the same size but different random seeds, providing further evidence for the transferability of DMs.333 The expected overlap between two random DMs is 0.33.
 4.2 Transferability Across Pretraining Methods
+We now study the transferability of training dynamics across different pretraining methods.
+Successful transfers of this type enable the use of efficient reference models, when there is no efficient version of the main model that suits the downstream task.
+The results for as the main model with different reference models are shown in Table 1 under section across different pretraining methods (NLI) and Table 5 (HSD, Appendix B).
+Training dynamics are generally transferable across different pretraining methods: achieves comparable performance using DMs constructed by different reference models in most cases.
+However, there is one exception: when using as the reference model, the performance is clearly worse on the NLI OOD datasets than when using ERM.
+We hypothesize that is not strong enough for constructing effective DMs.
+We analyze this hypothesis further in §4.3 below.
+4.3 When Do Transfers Fail?
 Method Main Model Ref.
-Model Cost MultiNLI AdversarialNLI (Test) - R1 R2 R3 Baselines: ERM, ERM with Early Stopping, and Random DM ERM - 14.47 ERM - 28.29 ERM* - 100.00 ERM(ES) - 66.67 DM Random 100.00 Training Dynamics Transferability: Across Different Model Sizes DM 200.00 DM 114.47 DM 128.29 Training Dynamics Transferability: Across Different Pretraining Methods DM 104.61 DM 136.18 DM 216.78 DM 213.49 FTFT: Efficient Reference Models + Aggressive Early Stopping FTFT 51.97 FTFT 74.12 FTFT 79.93
-We now study the transferability of training dynamics across different pretraining methods.
-Successful transfers of this type enable the exploitation of efficient reference models, in cases where there is no efficient version of the main model that suits the downstream task.
-The results for as the main model with different reference models are shown in Table 1 under section across different pretraining methods (NLI) and Table 5 (HSD, Appendix B).
-We observe that training dynamics are generally transferable across different pretraining methods: achieves comparable performance using DMs constructed by different reference models in most cases.
-However, there are a few exceptions: when using as the reference model, the performance is clearly worse on the NLI OOD datasets than when using ERM.
-We hypothesize that is not strong enough for constructing effective DMs, and analyze this hypothesis further in §4.3 below.
-4.3 How Efficient Can We Be?
-Method Main Model Ref.
-Model Compute MultiNLI (Val.) AdversarialNLI (Test) - R1 R2 R3 ERM TinyBERT - 1.45 ERM - 4.61 ERM - 36.18 ERM - 14.47 ERM - 28.29 ERM - 113.49 ERM - 116.78 DM TinyBERT 101.45 DM 104.61 DM 136.18 DM 114.47 DM 128.29 DM 213.49 DM 216.78 DM TinyBERT 111.64 DM 114.80 DM 146.38 DM 124.67 DM 138.49 DM 223.68 DM 226.97
+Model Compute MultiNLI (Val.) AdversarialNLI (Test) - R1 R2 R3 ERM Performance: Capabilities of Various Models ERM TinyBERT - 1.45 ERM - 4.61 ERM - 36.18 ERM - 14.47 ERM - 28.29 ERM - 113.49 ERM - 116.78 When Do Transfers Fail: Using Reference Models with Different Capabilities DM TinyBERT 101.45 DM 104.61 DM 136.18 DM 114.47 DM 128.29 DM 213.49 DM 216.78 DM TinyBERT 111.64 DM 114.80 DM 146.38 DM 124.67 DM 138.49 DM 223.68 DM 226.97
 We have shown that training dynamics are usually transferable across different model sizes and pretraining methods.
@@
 To answer this question, we compare the performance of a wide range of methods, see Table 2 (NLI) and Table 6 (HSD, in Appendix B).
-We include models from three categories: First, a wide range of seven models trained using ERM: TinyBERT, the small and base versions of DeBERTaV3 and ELECTRA, , and .
-Second, we use these ERM models as reference models to construct DMs, and use these DMs to fine-tune .
+We include models from three categories: First, a wide range of seven models trained using ERM (section ERM Performance): TinyBERT, the small and base versions of DeBERTaV3 and ELECTRA, , and .
+Second, we use these ERM models as reference models to construct DMs, and use these DMs to fine-tune (section When Do Transfers Fail, first half).
 By using reference models with different sizes and pretraining methods to fine-tune the same main model, we can inspect the impact of reference model capability on transferability.
-Third, we also include the results of using as the main model.
+Third, we also include the results of using as the main model (section When Do Transfers Fail, second half).
 By comparing results from different main models, we can better understand whether successful transfers originate from the compatibility between reference and main models, or the capability of the reference model itself.
@@
 First, weak reference models with poor ID performance, i.e., TinyBERT and , lead to unsuccessful transfers.
-However, the reference models do not need to be as capable as the main model: slightly weaker reference models could be more useful, indicated by the strong OOD performance of using as the reference model.
+However, the reference models do not need to be as capable as the main model: slightly weaker reference models could be more useful, indicated by the strong OOD performance when using as the reference model.
 Moreover, in these unsuccessful transfers, the main model OOD performance correlates with the reference model ID performance.
 For example, on NLI, TinyBERT has a lower ID performance than , and both main models and have a lower OOD performance when using TinyBERT as the reference model compared to when using as the reference model.
-Second, whether a transfer is successful or not depends mostly on the reference model, rather than the compatibility between the reference and the main models: The transfers from TinyBERT and to both main models are unsuccessful, and the transfers from other reference models to both main models are successful.
+Second, whether a transfer is successful or not depends mostly on the reference model, rather than the compatibility between the reference and the main models: The transfers from TinyBERT and to both main models are unsuccessful, and the transfers from other reference models to both main models are successful.444 We also include results using as the main model, and TinyBERT and as reference models on HSD in Table 6.
+Transfers from substantially outperform ERM, while those from TinyBERT underperform ERM, further suggesting that the reference model is decisive for successful transfers.
 Third, interestingly, ID performance in all transfers remains high.
@@
 We use median values because they are robust statistics of the central tendency.
-Figure 1(b) shows our results on MultiNLI, using two effective ( and ) and two ineffective ( and TinyBERT) reference models.
+Figure 2 shows our results on MultiNLI, using two effective ( and ) and two ineffective ( and TinyBERT) reference models.
 We only include four models to make the figure clearer — we observe similar trends for other models: With effective reference models, hard-to-learn instances are gradually learned during training, while other instances already have high values from the first epoch.
 In contrast, with ineffective reference models, hard-to-learn data instances are not learned at all, suggested by their close-to-zero over different epochs; while other instances are gradually fitted, indicated by their increasing values.
-To further validate our reasoning, we compute the percentages of training instances identified as easy by different reference models (Table 4 in Appendix B): We report results using to obtain more comprehensive results.
+To further validate our reasoning, we compute the percentages of training instances identified as easy by different reference models (Table 4 in Appendix B): We report results with .
 Less effective reference models indeed identify fewer data points as easy.
 For example, on NLI with , TinyBERT identifies less than 20.0% of the instances as easy, compared to 46.65% by .
-Furthermore, the overlap between hard-to-learn and ambiguous instances in successful transfers is usually very large.
-For example, with , all effective reference models identify more than 46% as easy training instances (the maximum is 50%, when ambiguous and hard-to-learn data align perfectly).
+Furthermore, the overlap between hard-to-learn and ambiguous instances in successful transfers is usually high.
+For example, with , all effective reference models identify more than 46% as easy training instances (the maximum is 50%, when the ambiguous and hard-to-learn subsets overlap perfectly).
 5 Training Speed Gains From Data Selection
-Our results from §4 suggest that one can improve the efficiency of dataset cartography with more efficient reference models.
-However, if we train the main models for the same number of steps as ERM, the cost of the full pipeline is still higher than conventional fine-tuning, because of the extra reference model training cost.
-Previous studies have shown that training with a careful selection of “informative” training instances can lead to higher training speed than ERM (i.e.
-achieving the same performance with fewer training steps), despite the computational cost of selecting these instances being notoriously high (Sorscher et al., 2022; Feldman & Zhang, 2020; Paul et al., 2021; Toneva et al., 2019).
+Our results from §4 suggest that the efficiency of dataset cartography can be improved with more efficient reference models.
+However, if we train the main models for the same number of steps as ERM, the computational cost of the full pipeline is still higher than conventional fine-tuning (i.e., ERM), because of the extra reference model training cost.
+Previous studies have shown that training with a careful selection of “informative" training instances can lead to higher training speed than ERM (i.e.
+achieving the same performance with fewer training steps), despite the computational cost of selecting these instances being notoriously high (Sorscher et al., 2022; Feldman and Zhang, 2020; Paul et al., 2021; Toneva et al., 2019).
 Motivated by this, we now study whether we can achieve similar learning speed gains by training on instances selected by DMs.
-If such gains exist, we can further improve the efficiency of data cartography by training with fewer steps.
-We present the performance of on HSD, fine-tuned using ERM and training instances selected by DMs, across different training durations (i.e.
+If such gains exist, we can further improve the efficiency of dataset cartography by training with fewer steps.
+We show the performance of on HSD, fine-tuned using the full training set (ERM) and using only the instances selected by DMs, across different training durations (i.e.
 max training steps) in Figure 1.
-We include results for other datasets in Appendix B, on which we observe similar trends.
-Specifically, we make three observations.
-First, when the number of training steps is reduced, models fine-tuned with dataset cartography clearly outperform those fine-tuned with ERM, even in cases where ERM achieves better performance with more training steps (i.e., on ID validation data).
-Second, dataset cartography often yields better results with fewer training steps than with more, suggesting a tendency of overfitting in the late stage of training.
-Third, such efficiency gains hold consistently across different training durations.
-All three observations indicate that training with data instances selected by DMs achieves consistent higher training speed than ERM.
+Results on other datasets are in Appendix B, where we observe similar trends.
+We make three observations.
+First, when the number of training steps is reduced, models fine-tuned with dataset cartography clearly outperform those fine-tuned with ERM, even on ID validation data where ERM achieves better performance with more training steps.
+Second, dataset cartography often yields better results with fewer training steps than with more, suggesting a tendency of overfitting at the late training stage.
+Third, the efficiency gains hold consistently across different training durations.
+Our observations indicate that training with data instances selected by DMs consistently achieves higher training speed than ERM.
 6 Our Approach: FTFT
-Building on our insights presented in the previous sections, we now propose a novel fine-tuning approach based on dataset cartography, Fine-Tuning by transFerring Training dynamics (FTFT).
-Compared with the original dataset cartography approach, FTFT integrates two crucial changes: (1) using more efficient reference models, which not only result in comparable DMs but also enhance efficiency, and (2) implementing aggressive early stopping during the main model training.
+Building on our insights presented in the previous sections, we propose a novel fine-tuning approach based on dataset cartography (Swayamdipta et al., 2020), Fine-Tuning by transFerring Training dynamics (FTFT).
+Compared with the original dataset cartography approach, FTFT integrates two crucial improvements: (1) using more efficient reference models, which not only result in comparable DMs but also enhance efficiency, and (2) implementing aggressive early stopping during the main model training.
 Such aggressive early stopping is possible because models trained with data instances selected using DMs already achieve strong performance with substantially fewer training steps (§5).
+In summary, FTFT consists of three steps: (1) train an efficient reference model on the full dataset, (2) compute the ambiguity of each instance based on the standard deviation of the correct class probability across epochs, and (3) select the top most ambiguous instances to retrain a larger main model, while applying aggressive early stopping.
 We show our results on NLI in Table 1 (FTFT) and on HSD in Table 5 (Appendix B).
 We use an early stopping patience (i.e.
-the number of checkpoints without improvement before stopping), based on the average dev set performance of the OOD datasets.
-We use a relatively small , because it can stop the training earlier, leading to higher efficiency.
+the number of checkpoints without improvement before stopping), based on the average dev set performance on the OOD datasets.
+We use a relatively small , as stopping the training earlier will lead to higher efficiency.
 We also show the computational cost of each method (Compute), in which we have taken the training of both the reference model and the main model into account, including the extra checkpoints after the best performance is achieved.
-We make three observations.
-First, FTFT achieves consistent robustness improvement over ERM, indicated by its strong performance on most OOD datasets (note again that the hyper-parameters are optimized for ERM).
+We have three key observations.
+First, FTFT achieves consistent robustness improvements over ERM, indicated by its strong performance on most OOD datasets (note again that the hyper-parameters are optimized for ERM).
 We also include the results of ERM with early stopping (ERM(ES)) as a baseline to illustrate that our aggressive early stopping strategy is only effective when training with data instances selected by DMs: ERM(ES) achieves worse performance than ERM on all OOD datasets.
-Second, FTFT achieves substantial higher training efficiency than other methods.
-For example, on NLI, training using FTFT and as reference models only costs 51.57% of the training cost of ERM, which is equivalent to 25.78% of the training cost of the original dataset cartography approach.
+Second, FTFT substantially improves the training efficiency of models.
+For example, on NLI, training using FTFT and as the reference model only costs 51.57% of ERM’s training cost, which is equivalent to 25.78% of the training cost of the original dataset cartography approach.
 Third, our early stopping strategy is effectively aggressive.
-Specifically, FTFT with all three reference models in Table 1 achieves their best performance within just 1/3 of the optimal training steps of ERM.
-7 Conclusions
-Fine-tuned PLMs have been shown to be vulnerable to OOD inputs.
-Although dataset cartography can improve model robustness (Swayamdipta et al., 2020; Sar-Shalom & Schwartz, 2023), it is computationally expensive.
+Specifically, all FTFT results in Table 1 are achieved using less than 1/3 of the optimal training steps needed for ERM.
+7 Conclusion
+Fine-tuned PLMs have shown to be vulnerable to OOD input.
+Although dataset cartography can improve model robustness (Swayamdipta et al., 2020; Sar-Shalom and Schwartz, 2023), it is computationally expensive.
 In this paper, we have presented FTFT, a novel approach for fine-tuning PLMs which yields both better efficiency and better robustness over ERM (§6).
-FTFT is built on dataset cartography, based on two observations: (1) reference model training dynamics are highly transferable across different model sizes (§4.1) and pretraining methods (§4.2), and (2) main models trained using data instances selected by reference model training dynamics achieve higher learning speed (i.e.
-substantially better performance with reduced number of training steps).
-We have also discussed the conditions for successful FTFT runs (§4.3).
+FTFT is built on dataset cartography, based on two observations: (1) reference model training dynamics are highly transferable across different model sizes (§4.1) and pretraining methods (§4.2), and (2) main models trained using data instances selected using DMs learn faster (i.e.
+they have substantially better performance with reduced number of training steps).
 We believe that FTFT will be an important tool for future researchers and practitioners to perform efficient model fine-tuning, especially in situations where robustness is essential.
-We also believe our study of lightweight reference models will benefit future research on building efficient supervised data pruning metrics, which are shown to outperform their unsupervised counterparts (Sorscher et al., 2022).
-Our work opens up several directions for future research.
+8 Limitations
+We identify three limitations from this work.
+These limitations open up promising directions for future research.
 First, we have observed that effective reference models identify more instances as easy.
-Further controlled experiments are needed to build a more rigorous protocol for choosing reference models.
-We believe the results presented here provide a good foundation for such further work.
-Second, we have empirically identified the transferability of training dynamics to selected training instances.
+More controlled experiments are needed to build a detailed protocol for choosing efficient but strong enough reference models.
+Our results provide a good foundation for such further work.
+Second, we have empirically demonstrated the transferability of training dynamics to select training instances.
 Future studies are needed to build the theoretical foundations of both data cartography itself and the feasibility of such transfers.
 Third, we have only developed FTFT for classification tasks.
-Future studies can extend our work to other tasks, such as generation (e.g., question answering) and self-supervised tasks (e.g., language modeling).
+Future studies should extend FTFT to generation tasks, e.g., instruction following and language modeling.
+Acknowledgments
+We thank the anonymous reviewers for their helpful feedback. This work used the Dutch national e-infrastructure with the support of the SURF Cooperative using grants no. EINF-10999 and EINF-5693.
 Appendix A Training Specifications
@@
 Optimization
-For training all models, we use AdamW (Loshchilov & Hutter, 2019) as the optimizer with a batch size of 32.
+For training all models, we use AdamW (Loshchilov and Hutter, 2019) as the optimizer with a batch size of 32.
 We also use a linear learning rate scheduler with 10% warmup.
 For fine-tuning the small and base versions of both DeBERTaV3 and ELECTRA, as well as TinyBERT, RoBERTa, and BERT, we use a learning rate of 2e-5, following the suggestions from the original papers.
-We use a smaller learning rate of 1e-5 for and , because training larger models with lower learning rates are observed to be more stable (Mosbach et al., 2021; Du & Nguyen, 2023).
-However, a few failed training runs (i.e., training fails to converge and the resulting model performs worse than the majority class baseline Mosbach et al., 2021) are still observed during the training of .
+We use a smaller learning rate of 1e-5 for and , because training larger models with lower learning rates are observed to be more stable (Mosbach et al., 2021; Du and Nguyen, 2023).
+However, a few failed training runs (i.e., runs where the training fails to converge and where the resulting model performs worse than the majority class baseline (Mosbach et al., 2021)) still occurred during the training of .
 We excluded these runs from our results.
@@
 | 4.40M | 1.45 |
+We estimate the training FLOPs by their models sizes (i.e.
+the number of parameters), following Kaplan et al.
+(2020) (, where is the number of FLOPs, is the number of parameters, and is the dataset size).
+The results are in Table 3.
+We prefer theoretical estimation over practical measurement, because (1) the results are less influenced by noises, (2) it considers both forward and backward propagation, and (3) it has been shown effective and widely used (Kaplan et al., 2020; Hoffmann et al., 2022).
+In practice, we also measured the actual FLOPs usage in a forward run using Sovrasov (2018), and the results are very close to our estimation using model sizes.
 Appendix B Additional Results
@@
 | - | Ori-R2 | Ori-R3 | Ori-R4 | Pert-R2 | Pert-R3 | Pert-R4 | ||||
+| ERM Performance: Capabilities of Various Models | ||||||||||
 | ERM | TinyBERT | - | 1.45 | |||||||
@@
 | ERM | - | 116.78 | ||||||||
+| When Do Transfers Fail: Using Reference Models of Different Capabilities | ||||||||||
 | DM | TinyBERT | 101.45 | ||||||||
@@
 | DM | 226.97 | |||||||||
+| DM | TinyBERT | 29.74 | ||||||||
+| DM | 42.76 | |||||||||